A sample of argon gas at a pressure of 0.894 atm and a temperature of 199 °C, occupies a volume of 853 mL. If the gas is heated at constant pressure until its volume is 998 mL, the temperature of the gas sample will be .... °C.
Answer #1

Charles law states that volume of a gas is directly proportional to its absolute temperature at constant pressure.

V1/T1 = V2/T2

T2 = V2 T1/V1 = 998 mL x (199+273) K / 853 mL = 552 K

Temperature of gas = 552 - 273 = 279 oC